OpenStudy (anonymous):

Marvin was completing the square, and his work is shown below. Identify the line where he made his mistake. f(x) = 2x^2 − 8x + 5 Line 1: f(x) = 2(x^2 − 8x) + 5 Line 2: f(x) = 2x^2 − 8x + 16) + 5 − 32 Line 3: f(x) = 2(x − 4)^2 − 27

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} \qquad f(x) = 2x^2 - 8x + 5\\ Line 1: f(x) = {\color{blue}{ 2}}(x^2 - {\color{blue}{ 8}}x) + 5\\ Line 2: f(x) = 2x^2 - 8x + 16) + 5 - 32 \\ Line 3: f(x) = 2(x - 4)^2 − 27 \end{array}\) what do you think?